OpenStudy (anonymous):

Find the limit for lim (sin4x)/(sin6x) x==>0

OpenStudy (helder_edwin):

gotta use \[ \large \lim_{x\to0}\frac{\sin x}{x}=1 \]

OpenStudy (anonymous):

But how do I use that when I have 4x and 6x?

OpenStudy (anonymous):

What are the steps to solving this? \[\lim_{x \rightarrow 0} \tan(6x)\div \sin(2x)\]

OpenStudy (anonymous):

So I can find the limit

OpenStudy (helder_edwin):

sorry. busy

OpenStudy (helder_edwin):

first multiply and divide by 1/x \[ \large \lim_{x\to0}\frac{\sin 4x}{\sin 6x}= \lim_{x\to0}\frac{\frac{\sin 4x}{x}}{\frac{\sin 6x}{x}} \]

OpenStudy (helder_edwin):

then multiply and divide the numerator by 4 and by 6 the denominator